# Analytics, have some humility: a statistical view of fourth-down decision making

### 11 Dec 2023 v2

@@ Abstract
The standard mathematical approach to fourth-down decision making in American football is to make the decision that maximizes estimated win probability. Win probability estimates arise from a statistical model fit from historical data. These machine learning models, however, are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. We develop a machine learning framework that accounts for this auto-correlation and knits uncertainty quantification into our decision making. In particular, we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. Our final product is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim.
@@ 1 Introduction
The two most widely used value functions by analysts of American football are expected points and win probability. Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next scoring event in the half given the current game-state. Net points is relative to the team with possession (negative if the opposing team is expected to score more points next). The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Initial solutions from Romer, (2006) and Burke, (2009) suggest making the decision that maximizes expected points. A team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function and ${\mathsf{WP}}$ is the right objective function. Hence modern solutions from Baldwin, (2021) and Burke suggest making the decision that maximizes win probability.111 Burke’s method is proprietary. Burke (Twitter: @bburkeESPN) and Baldwin (Twitter: @ben_bot_baldwin) post their fourth down recommendations on Twitter during and after real football games. These analyses found that NFL coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
Win probability estimates arise broadly from one of two classes of models, probabilistic state-space models or statistical models. State-space models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and are then propagated into win probability by simulating games. When done right, these models are a good way to estimate ${\mathsf{WP}}$, but they are difficult to do: they require a careful encoding of the convoluted rules of football into a set of states and the actions between those states, careful estimation of transition probabilities, and enough computing power to run enough simulated games to acheive desired granularity. Each of these are nontrivial.
On the other hand, statistical models are fit entirely from historical data. Given the results of a set of observed football plays, statistical models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. These models are widely used today in the football analytics community because rich publicly available play-by-play data (e.g., nflFastR (Carl and Baldwin,, 2022)) and powerful off-the-shelf machine learning models (e.g., ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016)) have become widely accessible. Additionally, due to a perceived abundance of data, these machine learning models are viewed as more trustworthy than previous mathematical models that make more simplifying assumptions. For these reasons, we focus on statistical win probability models in this paper.
With today’s abundance of big data and machine learning models, the typical approach to data-driven decision making is: fit a machine learning model to observational data and use the resulting point estimates to make decisions. This is the approach taken by the football analytics community: use machine learning to fit a win probability model from historical data and make the decision that maximizes win probability according to point estimates from this model. Machine learning models fit from noisy or small datasets, however, can overfit and have high variance. Trusting point estimates from these models can then lead to suboptimal decision making. This is particularly germane to today’s typical mathematical approach to fourth down decision making.
@@
In particular, the historical dataset of all football plays in recent history consists of highly correlated and noisy observations. Concisely, every game has only one winner. The binary win/loss response values are not independent, as all plays from the same game share the exact same draw of the win/loss outcome. The correlated nature of football play-by-play data reduces the effective sample size, exacerbating overfitting and inflating ${\mathsf{WP}}$ standard errors. Existing statistical win probability models do not account for this correlation and produce high variance estimates. This uncertainty should percolate into the fourth-down decision procedure, which currently treats ${\mathsf{EP}}$ and ${\mathsf{WP}}$ as known quantities rather than estimates. Trusting the resulting ${\mathsf{WP}}$ point estimates to dictate fourth-down decision making without accounting for uncertainty in these estimates leads to overconfident and potentially suboptimal decisions.
@@
Hence in this study we use bootstrapping to quantify uncertainty in fourth-down recommendations. Our confidence in a decision recommended by a win probability point estimate is proportional to the number of bootstrapped models that report that decision as optimal. In other words, we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. Our final product is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim. A substantial proportion of fourth-down decisions feature win probability recommendations subject to large uncertainty. Thus, we ask football analysts to have some humility: for many game-states, there is simply not enough data to know which decision is optimal.
@@ 2 The traditional fourth-down decision procedure
The standard mathematical approach to fourth-down decision making is to make the decision that maximizes estimated win probability. In this Section we explore the traditional decision procedure and detail how analysts typically estimate win probability. We begin with a brief overview of our historical dataset of football plays in Section 2.1. Then in Section 2.2 we illustrate the traditional decision process for example plays. In Section 2.3 we describe how to estimate a team’s fourth-down win probability if it punts, kicks a field goal, or attempts a conversion. Fourth-down win probability is typically estimated in terms of first-down win probability, so in Section 2.4 we describe well known open source first-down win probability models.
@@ 2.1 Data
We scraped every football play from 1999 to 2022 from the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes over 200 variables such as score differential, total score, game seconds remaining, half seconds remaining, yardline, down, yards to go, whether the team with possession is at home, whether the team with possession receives the second half kickoff, the number of timeouts remaining for each team, whether the team with possession wins the game, and other variables relevant to estimating win probability. Note that yardline is a real number in $\{0,1,...,99,100\}$ measuring the number of yards from the opponent’s endzone, where 00 represents a touchdown and $100$ indicates a safety. We hold out plays from 1999 to 2005 to avoid data bleed in constructing kicker quality, punter quality, and team quality metrics in Appendix B. We hold out plays from 2022 to illustrate our decision procedure on out-of-sample example plays in Section 4.1. We are left with a primary dataset of $600,825$ plays, with $240,380$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models from this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.222 https://github.com/snoopryan123/fourth_down
@@ 2.2 Example plays
The traditional decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. Further, the strength of a decision is proportional to the estimated gain in win probability by making that decision. Notably, Baldwin, (2021) and Burke employ this decision procedure and post their fourth-down recommendations on Twitter.333 Baldwin’s posts his fourth-down recommendations at @ben_bot_baldwin and Burke posts his at @bburkeESPN. We detailed Baldwin’s first-down win probability model in Section 2.4. Burke’s model is proprietary. We illustrate this procedure through two example plays.
Example play 1. First we give an example of Baldwin’s fourth-down decision charts. Baldwin suggests making the decision that maximizes estimated win probability. Figure 1 illustrates Baldwin’s decision making charts for a fourth-down play in which the Colts had the ball against the Patriots in Week 10 of 2023. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
@@
Example play 2. Next we give an example of Burke’s decision chart. Burke, whose win probability model is proprietary, also suggests making the decision that maxizimes estimated win probability. Figure 2 illustrates Burke’s fourth-down decision chart for a play from the 2023 NFC Championship game. The chart visualizes the estimated optimal decision according to effect size (color) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary.
@@ 2.3 Estimating the win probability of a fourth-down decision
To make the decision that maximizes estimated win probability, we need to estimate a team’s fourth-down win probability if it punts (${\mathsf{Punt}}$), kicks a field goal (${\mathsf{FG}}$), or attempts a conversion (${\mathsf{Go}}$) as a function of game-state, which we describe in this Section.
To estimate win probability on fourth down, we could gather all the fourth-down plays in our dataset and learn whether ${\mathsf{Go}}$, ${\mathsf{FG}}$, or ${\mathsf{Punt}}$ leads to better game outcomes. Traditional fourth-down decision procedures do not take this approach due to a lack of data. In our dataset there are $94,786$ fourth downs and just $18\%$ of these ($17,196$ plays) are conversion attempts. Most of these conversion attempts appear late in games when a team obviously has to go for it. Few of these conversion attempts appear earlier in games or in moments when ${\mathsf{Go}}$ is plausibly but not obviously the best decision. So, estimating win probability if a team goes for it on fourth-down just from game outcomes won’t produce powerful estimates.
To gain power, we instead estimate fourth-down win probability in terms of first-down win probability and a decision transition probability. We can do this because a first down always follows a fourth down. For instance, success on a conversion attempt results in a first down for the team with possession and failure results in a first down for the other team. Thus ${\mathsf{Go}}$ win probability on fourth down is a function of conversion success probability and first-down win probability. As there are many more first-down plays in our dataset than fourth-down plays, it is easier to estimate first-down win probability than fourth-down win probability. Also, it is relatively straightforward to estimate decision transition probabilities (e.g., field goal success probability) because they are fit from i.i.d. play-level outcomes. Hence we gain power from estimating fourth-down win probability in terms of first-down win probability.
Although estimating fourth-down win probability in terms of first-down win probability produces more powerful estimates, it is still very difficult to accurately estimate first-down win probability from historical data. Our dataset consists of many first-down plays ($240,380$), but these plays come from relatively few games ($4,101$). Thus first-down win probability models, which are fit from game-level outcomes, are subject to a small effective sample size. In contrast, decision transition probabilities are fit from i.i.d. play-level outcomes and so are fit from a much larger effective sample size. On this view, estimating first-down win probability is the most difficult component of fourth-down decision making.
We focus on the crux of the fourth-down decision framework, first-down win probability, in the upcoming Sections. We relegate the remainder of fourth-down win probability modeling to the Appendix. Specifically, in Appendix A we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We estimate these decision transition probabilities in Appendix B.
@@ 2.4 Estimating the win probability of a first down
To produce better estimates, we compute fourth-down win probability in terms of first-down win probability. In this Section we describe well known open source win probability models, whose first-down estimates form the crux of the fourth-down decision procedure. These are statistical models fit from the historical dataset of football plays.
Baldwin, (2021) uses ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016) to estimate win probability from historical data. The response variable is a binary variable indicating whether the team with possession wins the game. Baldwin models win probability as a function of score differential, game seconds remaining, half seconds remaining, yardline, down, yards to go, whether the team with possession is at home, whether the team with possession receives the second half kickoff, and the number of timeouts remaining for each team. He uses two additional features to capture the change in impact of point spread and score differential over the course of a game,
@@
Baldwin includes monotonic constraints for yardline, yards to go, down, score differential, timeouts remaining for each team, spread-time, and diff-time-ratio. He tunes the ${\mathsf{XGBoost}}$ hyperparameters by minmizing cross validated log-loss (Baldwin,, 2021).
Lock and Nettleton, (2014) use a Random Forest (${\mathsf{RF}}$) (Breiman,, 2001) to estimate win probability from historical data. They use the same binary win/loss response variable as before. They model win probability as a function of score differential, game seconds remaining, yardline, down, yards to go, the number of timeouts remaining for each team, pre-game point spread, total points scored, and
adjusted score=score differential1+game seconds $\text{adjusted score}=\frac{\text{score differential}}{\sqrt{1+\text{game seconds remaining}}}.$ (2.3)
@@
Each of these win probability models are statistical models fit from a historical play-by-play dataset. This dataset includes plays from all downs; these models fit first-down win probability by including down as a covariate. We find that a model fit from just first-down plays better predicts first-down win/loss, which is what we care about for fourth-down decision making (see Appendix C). For the decision procedure used in the remainder of this paper, we use this first-down ${\mathsf{XGBoost}}$ win probability model fit from just first downs, described in Appendix C.
@@ 3 Fourth-down decision making under uncertainty
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. This procedure ignores the uncertainty inherent in estimating win probability from noisy and highly correlated historical data. Thus in this Section we modify the decision procedure to account for uncertainty. In Section 3.1 we introduce bootstrap percentage as a way to quantify uncertainty in the estimated optimal fourth-down decision. Specifically, we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. Then, to understand just how difficult it is to accurately fit a win probability model from historical data, in Section 3.2 we conduct a simulation study using a simplified random walk version of football in which the true win probability is known. Although statistical win probability models find the general trend of true win probability, they are subject to substantial uncertainty. Finally, in Section 3.3 we implement our full fourth-down decision procedure based on bootstrapping.
3.1 Quantifying uncertainty in the estimated optimal decision
The traditional decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. This procedure treats win probability estimates as known quantities, ignoring the fact that win probability models produce estimates. This begs the question: when can we trust the decision recommended by win probability point estimates?
@@ 3.1 Quantifying uncertainty in the estimated optimal decision
The open source first-down win probability models that form the crux of traditional fourth-down decision procedures today are machine learning models fit from historical data. Football analysts today typically use these models because they see a massive dataset – $600,825$ plays with $240,380$ first-down plays in our play-by-play dataset – and think this is enough data to fit a highly accurate model using machine learning. This, however, is not true because the binary win/loss response variable is noisy and highly correlated: every game has only one winner. Formally, the binary response variable $y_{i}$ of the $i^{th}$ play indicates whether the team with possession won the game. The reponse values are not independent, as all plays from the same game share the same draw of the response column. Thus the effective sample size is somewhere between the number of plays ($600,825$) and the number of non-tied games ($4,101$) in our dataset. This is likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yardline, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context, and as such we expect win probability point estimates to have high variance.
Therefore it is imperative to incorporate uncertainty quantification into the fourth-down decision procedure. Since we estimate win probability using machine learning models, we need a non-parametric method to quantify uncertainty in these estimates. Bootstrapping is a natural choice to capture such uncertainty.
The bootstrapping process begins with generating $B$ bootstrapped datasets from our original dataset. We describe the technical details of how we generate each bootstrapped dataset in Section 3.3 and how we choose an appropriate value for $B$ in Appendix E; these technical details are not necessary to understand the idea behind bootstrapping win probability. Each bootstrapped dataset mimics a new draw of the training set: keeping all of the pre-game attributes of each game fixed (e.g., each team’s composition), bootstrapping mimics re-drawing plays and the random win/loss outcome of each game. In other words, each bootstrapped dataset represents a re-draw of the recent history of football. Then, to each bootstrapped dataset we fit a fourth-down win probability model. At each game-state, each model yields an estimated optimal decision. The strength of each estimated optimal decision is proportional to the effect size, or the estimated gain in win probability by making that decision. Because the plays and win/loss outcomes in each bootstrapped dataset differ from those in the original dataset, each bootstrapped decision function differs from the original one.
There are two problems with using these confidence intervals to drive decision making. First, since a confidence interval is defined by two numbers, it is difficult to meld those two numbers into one decision recommendation. For example, consider a play in which the estimated gain in win probability for ${\mathsf{Go}}$ is $1\%$. If our confidence interval for this gain is $[-3\%,5\%]$, it is not clear how strongly we should recommend ${\mathsf{Go}}$. Similarly, it is difficult to compare two confidence intervals of varying lengths. For example, it is not clear whether our recommendation for a confidence interval of $[-3\%,5\%]$ should be weaker than that for a confidence interval of $[-1\%,3\%]$. Second, we are primarily interested in quantifying uncertainty in the decision itself (Friedman et al.,, 1999). We do not necessarily need to accomplish this by quantifying uncertainty in the effect size, which is what these confidence intervals do.
To boil a bootstrapped confidence interval on the effect size into a single summary statistic to drive decision making, an analyst may be tempted to average the $B$ bootstrapped effect sizes. Similarly, she may be tempted to consider the distribution of bootstrapped effect sizes or to integrate over them in some way. It is wrong to do this because bootstrapping does not produce a posterior distribution. The bootstrapped values $\hat{g}_{1}({\mathbf{x}}),...,\hat{g}_{B}({\mathbf{x}})$ are not guaranteed to follow any particular distribution and it does not make sense to integrate over these values.
Rather, to quantify uncertainty in the fourth-down decision itself we use boostrap percentage (${\mathsf{boot\%}}$), or the percentage of bootstrapped models that report each decision to be optimal. Formally, at game-state ${\mathbf{x}}$ we have the original estimated optimal decision $d({\mathbf{x}})\in\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ (the point estimate) and $B$ bootstrapped estimated optimal decisions $d_{1}({\mathbf{x}}),...,d_{B}({\mathbf{x}})$. Then
${\mathsf{boot\%}}({\mathbf{x}})=100\%\cdot\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\left(d_{b}({\mathbf{x}})=d({\mathbf{x}})\right).$ (3.1)
Bootstrap percentage quantifies how confident we are that a fourth-down recommendation has higher win probability than all the other decisions. A high bootstrap percentage (${\mathsf{boot\%}}\approx 100\%$) reflects high confidence in the estimated optimal decision: the estimated optimal decision remains the same across the vast majority of re-draws of the training set. When ${\mathsf{boot\%}}$ is high, we can trust the decision recommended by the point estimate. Conversely, a low bootstrap percentage reflects low confidence in the estimated optimal decision: the estimated optimal decision differs substantially across different re-draws of the training set. A low bootstrap percentage is ${\mathsf{boot\%}}\approx 50\%$ or lower because a fourth-down decision almost always boils down to picking between ${\mathsf{Go}}$ and one of the kicks, as it is clearly suboptimal to punt near the opponent’s endzone and to attempt a field goal far from the opponent’s endzone. Regardless of the effect size, when ${\mathsf{boot\%}}$ is low we cannot trust the point estimate: the effect size at that game-state is too dependent on the random idiosyncracies of its particular training set. In other words, the edge detected by the point estimate is more due to noise than signal. In a “parallel universe” in which we re-drew each random win/loss game outcome, the estimated optimal decision at that game-state may differ solely due to the randomness inherent in drawing the history of football.

### 19 Jan 2024 v4

@@ Abstract
The standard mathematical approach to fourth-down decision making in American football is to make the decision that maximizes estimated win probability. Win probability estimates arise from machine learning models fit from historical data. These models, however, are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. Thus, it is imperative to knit uncertainty quantification into the fourth-down decision procedure; we do so using bootstrapping. We find that uncertainty in the estimated optimal fourth-down decision is far greater than that currently expressed by sports analysts in popular sports media. Our contribution is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim.
Keywords: Decision making under uncertainty, applications and case studies, machine learning, bootstrap/resampling, statistics in sports
@@ 1 Introduction
The two most widely used value functions by analysts of American football are win probability and expected points (Yurko et al.,, 2018). Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next score in the game, relative to the team with possession, given the current game-state. The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Previous work by Romer, (2006) and Burke, 2009b suggest making the decision that maximizes estimated ${\mathsf{EP}}$. Yet a team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function. Hence, modern approaches by Baldwin, (2021) and Burke, 2009a suggest making the decision that maximizes estimated ${\mathsf{WP}}$. These analyses found that National Football League (NFL) coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
@@
The ${\mathsf{WP}}$ estimates used for fourth-down decision making typically arise from statistical models fit from historical data. Given the results of the entire recent play-by-play history of football, these models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. However, these models are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. Because every game has only one winner, the binary win/loss response values are not independent, as all plays from the same game share the exact same draw of the win/loss outcome. This reduces the effective sample size and inflates standard errors.
Traditional decision procedures, based solely on effect size, ignore the uncertainty inherent in estimating ${\mathsf{WP}}$ from noisy and highly correlated historical data. This uncertainty should percolate into the fourth-down decision procedure, but popular approaches currently treat win probabilities as known quantities rather than estimates. Trusting the resulting ${\mathsf{WP}}$ point estimates to dictate fourth-down decision making without accounting for uncertainty in these estimates leads to overconfident and potentially sub-optimal decisions.
@@
Our focus is not to “fix” win probability models by adjusting for additional covariates or reducing model bias. Rather, we shed light on the high variance nature of estimating win probability and show that such estimates are rife with uncertainty. We use bootstrapping to quantify uncertainty in fourth-down recommendations and recommend a fourth-down decision when we are confident it has higher ${\mathsf{WP}}$ than all other decisions. Our contribution to the literature is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim. A substantial proportion of fourth-down decisions feature ${\mathsf{WP}}$ recommendations subject to large uncertainty. Thus, we ask football analysts to have some humility: for many game-states, there is simply not enough data to trust ${\mathsf{WP}}$ point estimates.
@@ 2 The traditional fourth-down decision procedure
The standard mathematical approach to fourth-down decision making is to make the decision that maximizes estimated win probability. In this Section we explore the traditional decision procedure and detail how analysts commonly estimate win probability today. We begin with a brief overview of our historical dataset of football plays in Section 2.1. Then in Section 2.2 we illustrate the traditional decision process for example plays. In Section 2.3 we describe how analysts estimate a team’s fourth-down win probability if it punts, kicks a field goal, or attempts a conversion. Fourth-down win probability is typically estimated in terms of first-down win probability, so in Section 2.4 we describe well known open source first-down win probability models.
@@ 2.1 Data
We access every NFL play from 1999 to 2022 using the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes variables that are relevant to estimating win probability that describe the context of the play, such as the score differential, time remaining, yardline, down, yards to go, etc. (see Table 3 in Appendix A for descriptions of relevant variables). Note that yardline is a real number in $\{0,1,...,99,100\}$ measuring the number of yards from the opponent’s endzone, where 00 represents a touchdown and $100$ indicates a safety.
In order to avoid data bleed in constructing the various decision transition models (Appendix C) relevant for fourth-down decision making, we hold out plays from 1999 to 2005. Additionally, we hold out plays from 2022 to illustrate our decision procedure on out-of-sample example plays in Section 4.1. We are then left with a primary dataset of $600,825$ plays, with $240,380$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models using this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.111 https://github.com/snoopryan123/fourth_down
@@ 2.2 Example plays
The traditional fourth down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. Further, the strength of a decision is proportional to the estimated gain in win probability by making that decision. Notably, Baldwin, (2021) and Burke, 2009a employ this decision procedure and post their fourth-down recommendations on Twitter.222 Baldwin’s posts his fourth-down recommendations at @ben_bot_baldwin and Burke posts his at @bburkeESPN. We illustrate this procedure through two example plays.
Example play 1. First, we consider an example of Baldwin’s fourth-down decision charts, which is based on a model described in Section 2.4. Figure 1 illustrates Baldwin’s decision-making charts for a fourth-down play in which the Colts had the ball against the Patriots in Week 10 of 2023. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
@@
Example play 2. Next, we give an example of Burke’s decision chart, which is based on a proprietary model. Figure 2 illustrates Burke’s fourth-down decision chart for a play from the 2023 NFC Championship game. The chart visualizes the estimated optimal decision (color) according to effect size as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary.
@@ 2.3 Estimating the win probability of a fourth-down decision
An offense has three possible decisions to make on fourth down: it can punt the ball, kick a field goal, or attempt a conversion (colloquially known as “going for it”), denoted ${\mathsf{Punt}}$, ${\mathsf{FG}}$, and ${\mathsf{Go}}$, respectively. To employ the traditional fourth-down decision procedure, we need to estimate a team’s fourth-down win probability for each decision as a function of the game-state.
In Appendix B we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We can do this because a first down always follows a fourth down. For instance, success on a conversion attempt results in a first down for the team with possession and failure results in a first down for the other team. Thus, ${\mathsf{Go}}$ win probability on fourth down is a function of conversion success probability and first-down win probability. Similarly, ${\mathsf{FG}}$ win probability is a function of field goal success probability and first-down win probability. ${\mathsf{Punt}}$ win probability is a function of punt outcome and and first-down win probability. We estimate these decision transition probabilities in Appendix C.
We detail first-down win probability estimation in the main text, relegating the remainder of fourth-down win probability modeling to the Appendix, because it is the most difficult component of fourth-down decision making. Whereas decision transition models are fit from i.i.d. play-level outcomes and so have a relatively large effective sample size, first-down win probability models are fit from noisy, clustered, and highly correlated game-level outcomes. There are just $4,101$ games in our dataset, providing a relatively small effective sample size for statistical first-down win probability models. On this view, first-down win probability forms the crux of the fourth-down decision framework.
@@ 2.4 Estimating the win probability of a first down
Win probability estimates arise broadly from one of two classes of models, probabilistic state-space models or statistical models. On one hand, state-space models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and are then propagated into win probability by simulating games. When implemented correctly, these models are sensible ways to estimate ${\mathsf{WP}}$. However, they are difficult in practice, as they require: a careful encoding of the convoluted rules of football into a set of states and the actions between those states, careful estimation of transition probabilities, and enough computing power to run enough simulated games to achieve desired granularity. Each of these are nontrivial.
On the other hand, statistical models are fit entirely from historical data. Given the results of a set of observed football plays, statistical models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. These models are widely used today in the football analytics community because of publicly available play-by-play data (e.g., nflFastR (Carl and Baldwin,, 2022)) and accessible off-the-shelf machine learning models (e.g., ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016)). Additionally, due to a perceived abundance of data, flexible machine learning models are viewed as more “trustworthy” than previous mathematical models that make more simplifying assumptions. For these reasons, the open source win probability models used today for fourth-down decision making are statistical / machine learning models. Hence, in this paper we focus on statistical / machine learning win probability models.
Lock and Nettleton, (2014) use a Random Forest (${\mathsf{RF}}$) (Breiman,, 2001) to estimate win probability from historical data. The response variable is a binary variable indicating whether the team with possession wins the game. They model win probability as a function of score differential, game seconds remaining, yardline, down, yards to go, the number of timeouts remaining for each team, pre-game point spread, total points scored, and an additional feature to capture the change in impact of score differential over the course of a game,
adjusted score=score differential1+game seconds $\text{adjusted score}=\frac{\text{score differential}}{\sqrt{1+\text{game seconds remaining}}}.$ (1)
Baldwin, (2021) uses ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016) to estimate win probability from historical data. He uses the same binary win/loss response variable as before. Baldwin models win probability as a function of score differential, game seconds remaining, half seconds remaining, yardline, down, yards to go, whether the team with possession is at home, whether the team with possession receives the second half kickoff, and the number of timeouts remaining for each team. He uses two additional features to capture the change in impact of point spread and score differential over the course of a game,
@@
Baldwin includes monotonic constraints for yardline, yards to go, down, score differential, timeouts remaining for each team, spread-time, and diff-time-ratio. He tunes the ${\mathsf{XGBoost}}$ hyper-parameters by minimizing cross validated log-loss (Baldwin,, 2021).
@@
Each of these win probability models are statistical models fit from an historical play-by-play dataset. This dataset includes plays from all downs; these models fit first-down win probability by including down as a covariate. We find that a model fit from just first-down plays better predicts first-down win/loss (see Appendix D), which is what we care about for fourth-down decision making. For the decision procedure used in the remainder of this paper, we use a first-down ${\mathsf{XGBoost}}$ win probability model fit from just first downs as described in Appendix D.
@@ 3 Fourth-down decision making under uncertainty
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. The win probability models that form the crux of traditional fourth-down decision procedures today are machine learning models fit from historical data. Football analysts today typically use these models because they see a massive dataset – $600,825$ plays with $240,380$ first-down plays in our play-by-play dataset – and think this is enough data to fit a highly accurate model using machine learning. However, this is not true because the binary win/loss response variable is noisy and highly correlated: every game has only one winner. Formally, the binary response variable $y_{i}$ of the $i^{th}$ play indicates whether the team with possession won the game. The response values are not independent, as all plays from the same game share the same draw of the response column. Thus the effective sample size is somewhere between the number of plays ($600,825$) and the number of non-tied games ($4,101$) in our dataset. This is likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yardline, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context, and as such we expect win probability point estimates to have high variance.